OpenStudy (anonymous):

PLEASE HELP MEEE I HAVE TO GET THESE RIGHT OR ILL FAIL MATH THIS IS MY LAST CHANCE!!! Select the appropriate restrictions for the variables. 4a+13/a+3 A. a≠–3 B. a≠3 C. a≠4 D. There are no restrictions on the values of variable a.

OpenStudy (anonymous):

woah. hold on let me think i know this

OpenStudy (anonymous):

@johnweldon1993 please help me!!!

OpenStudy (anonymous):

hmmmmm i dont know sorry :(

OpenStudy (johnweldon1993):

Alright...so You have a fraction...The only thing you need to watch for with a fraction...is the denominator equaling out to 0 Why? You cannot divide a number by 0...it is undefined So what value for 'a' makes your denominator = 0?

OpenStudy (anonymous):

im confused

OpenStudy (johnweldon1993):

That's okay.. so we have \[\large \frac{4a + 13}{a + 3}\] right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Alright.. Let me put a visual with this... how does the graph of 1/x look?  Like that We can see from the graph...when x is any number...we an go up to the function line....but what about when x = 0? what happens we we go to 0 on the x-axis and try to go up or down to meet the function? We cannot do it...the function blows up and become undefined Does that make sense?

OpenStudy (anonymous):

no not at all

OpenStudy (anonymous):

omg ima cry im soo confused

OpenStudy (johnweldon1993):

Don't cry, We'll get you through it :)

OpenStudy (johnweldon1993):

Alright well basically like I've been saying....a denominator of a fraction can NEVER = 0 Because then you get a number divided by 0 ....which is undefined...we do not know it So with your question \[\large \frac{4a + 13}{\color \red{a + 3}}\] \(\large a + 3\) is your denominator....and since we know THAT can NOT equal 0...we do \[\large a + 3 = 0\] and now solve for 'a' What number...can you add to 3...to make 0?

OpenStudy (anonymous):

-3

OpenStudy (johnweldon1993):

Right....a can NOT = -3 because that will make your denominator = 0...and make this function undefined.... \(\large a \cancel{=} -3\)

OpenStudy (anonymous):

omg yay so i do that for a question like 5x-3x^2/2x^4

OpenStudy (anonymous):

right?

OpenStudy (johnweldon1993):

Is that... \[\large \frac{5x - 3x^2}{2x^4}\] right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Then yes exactly like that... What will make your denominator = 0? \[\large 2x^4 = 0\] What value of 'x' will make that true?

OpenStudy (anonymous):

-2x^4?

OpenStudy (johnweldon1993):

Not quite...remember this is multiplication hun :) it is 2 times x^4 \[\large 2 \times x^4 = 0 \] So 2 times what? = 0 ?

OpenStudy (anonymous):

2 times 0 =0

OpenStudy (johnweldon1993):

Right....so here \[\large x \cancel{=} 0\]

OpenStudy (anonymous):

okay..???

OpenStudy (johnweldon1993):

Alright, since your confused, let me ask is that question the same as this one? "state the restrictions on the variable 'x'?

OpenStudy (anonymous):

what does that mean omg lol

OpenStudy (johnweldon1993):

Lol okay okay.... Your first post on this... "Select the appropriate restrictions for the variables." And we found the restrictions to be 'a' can NOT equal -3 Is THIS the same type of problem? are we still "Select the appropriate restrictions for the variables." ?

OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variables. are the directions

OpenStudy (johnweldon1993):

Alright then yes we are doing the SAME thing as we did before ...okay? We are finding what value of 'x' we can plug in to make our new denominator = 0 Our new denominator is 2x^4 So 2x^4 = 0 if we plug in x = 0....we get 0....so our restriction is x can NOT = 0 That is your answer :)

OpenStudy (anonymous):

thats none of my answers though

OpenStudy (johnweldon1993):

Let me see your choices

OpenStudy (johnweldon1993):

Oh that's because you gave me a different question! :P you gave me \(\huge \frac{5x - 3x^2}{2x^4}\) But the question you are woking on is \(\huge \frac{(x+ 3)(x + 7)}{3x + 21} \)

OpenStudy (anonymous):

ohh i did im sorry